### v2

# Black hole event horizons are cosmologically coupled

Abstract
It is shown that an exactly static and spherically symmetric black hole event horizon cannot be embedded in a time-dependent geometry. Forcing it to do so results in a naked null singularity at the would-be horizon. Therefore, since the universe is expanding, black holes must couple to the cosmological expansion, which was suggested as the growth mechanism for supermassive black holes in galaxies, with implications for the dark energy puzzle.
I Introduction
A major problem of contemporary astrophysics consists of the fact that supermassive black holes at the centres of galaxies must have grown very fast at redshifts $z>6$, which is problematic if one invokes standard astrophysical channels [1, 2]. Moreover, LIGO detections of gravitational waves from binary mergers involving black holes highlight the presence of stellar mass black holes in the mass gap where they should not exist according to well-established models of stellar evolution [3, 4, 5], which is another potential problem if confirmed. These astrophysical puzzles can potentially be resolved if black holes interact with the universe in which they are embedded and their masses evolve over cosmological time scales (“cosmological coupling”) [6] 111 To be specific, in this paper “cosmological coupling” refers to any evolution of a black hole event horizon with time or with the scale factor of the FLRW in which they are embedded. .
These astrophysical black holes are usually modelled as asymptotically flat Schwarzschild or Kerr black holes. However, realistic black holes are not asymptotically flat but are embedded in the universe. The study of astrophysical black holes usually neglects the cosmic expansion due to the vast separation between astrophysical (local) scales and the cosmological scale. However, over billions of years, tiny interactions between black hole and surrounding universe, if present, accumulate and can become important. The recent Ref. [6] reports plausible evidence for cosmological coupling between the masses of supermassive black holes and the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe, obtained from data of black hole populations in red elliptical galaxies in the redshift range $0<z\leq 2.5$. Furthermore, if black holes are non-singular objects with de Sitter-like macroscopic cores, they could even explain dark energy, which would effectively be relegated to the interiors of black holes [6], a phenomenology possibly extending to non-singular black hole mimickers, which are now the subject of a vast theoretical literature (e.g., [7, 8, 9, 10, 11, 13, 14, 15, 16]. Cosmological coupling is also in agreement with the recent DESI results on the redshift evolution of the dark energy equation of state [17]. The 2023 tentative observation of cosmological coupling [6] caused a lively debate in both the observational and theoretical communities during the past year ([18, 19, 20, 21, 22] and references therein), as the arguments supporting it seemed unconvincing to some of its critics - here we provide more convincing ones.
Another context in which black holes can potentially interact with their cosmological surroundings is given by primordial black holes in the early universe, in which the Hubble scale is small and the separation between black hole and cosmological scales may not be sharp. Ref. [23] pointed out the inadequacy of the Schwarzschild (or Kerr) metric to describe primordial black holes. Although the analysis relied on the specific Thakurta solution of the Einstein equations that ultimately turns out not to describe a black hole [24, 25], its main point remains.
The problem of principle of describing a black hole or other strong field object immersed in a FLRW universe has a long history dating back to McVittie [26] and Einstein and Straus [27] (see [28] for a review of relevant solutions of the Einstein equations). Only recently, however, cosmological coupling (independent of accretion or mergers) has been included in realistic astrophysics ([6, 29] and references therein). In spite of the extremely interesting potential implications, the cosmological coupling of black holes does not yet have solid theoretical foundations, at least locally, nor it is backed by exact solutions of the Einstein equations possessing all the desirable properties for coupling, although progress was made in this direction [30, 31]. The possibility that black holes couple to the cosmological expansion, and that this coupling has even been observed, has triggered a debate in the community. Critics point to exact coupled solutions without any mass growth such as the Schwarzschild-de Sitter/Kottler solution (e.g., [33, 32]). A counter-argument is that, being general relativity a nonlinear theory, the solutions of its initial/boundary value problems are not always unique (e.g., [17]). To address this issue, here we turn around the question and ask whether an exactly static black hole event horizon (a null surface and a causal boundary) can be embedded in a FLRW universe. [^2] Because it is locally static, we exclude the de Sitter universe from our discussion (indeed, the Schwarzschild-de Sitter/Kottler solution is locally static and contains a black hole event horizon, but it is irrelevant for the problem addressed here). Interestingly, this question was raised in [34], when no one was yet taking cosmological coupling seriously. These authors assume a solution of the Einstein equations that is spherically symmetric, time-dependent, and is sourced by a perfect fluid with linear barotropic equation of state $P=w\rho$ (with $w=$ const.), answering the question negatively through a perturbative analysis [34]. Here we generalise significantly the result of [34] by studying the existence of an exactly static black hole horizon in a generic time-dependent geometry and provide several no-go arguments. In particular, we do not assume the Einstein equations nor specific forms of matter, but we rely only on a near-horizon expansion well-established in black hole physics. We find that embedding an exactly static black hole horizon in a time-dependent geometry turns it into a naked null singularity, providing theoretical support for cosmological coupling.
The next section sets up the description of an exactly static black hole event horizon in a time-dependent geometry, while the following sections show how this assumption leads to absurdities. Specifically, Sec. III, shows that null and timelike radial geodesics are incomplete at the would-be static horizon, where the Ricci scalar also diverges. This static horizon must coincide with a naked spacetime singularity at a finite (proper) radius. In Sec. IV we analyze the propagation of a test scalar field, which exhibits divergences at the would-be static horizon. Sec. V assumes a FLRW “background” and argues that the static would-be horizon leads to a paradox for the Hawking temperature, while Sec. VI contains the conclusions.
II Static black hole event horizon in a time-dependent metric
Let us set up the description of a static black hole event horizon embedded in a time-dependent “background” geometry which, remembering the origin of our problem, can be thought of as a FLRW universe. However, we do not make this assumption until Sec. V and the time-dependent metric is general, except for being spherically symmetric. We do not want to commit to any exact solution of the Einstein equations (or of the field equations of other theories of gravity) because these are ad hoc, and only a handful of them are known anyway [28]. In fact, we do not even assume the Einstein equations.
Without loss of generality, a spherically symmetric and time-dependent line element can be written in the form
$ds^{2}=-T^{2}(t,r)dt^{2}+a^{2}(t,r)\left(dr^{2}+r^{2}d\Omega_{(2)}^{2}\right)$ (2.1)
in isotropic coordinates. $d\Omega_{(2)}^{2}\equiv d\vartheta^{2}+\sin^{2}\vartheta d\varphi^{2}$ is the line element on the unit 2-sphere and the areal radius (a geometric quantity defined invariantly) is
$R(t,r)=a(t,r)r\,,$ (2.2)
as can be read off the angular part of the line element (2.1). The equation locating the apparent horizons [^3] This equation generalizes the well known relation locating the Schwarzschild event horizon $g^{RR}=\nabla^{c}R\nabla_{c}R=1-2m/R=0$ to any spherical geometry. In spherical symmetry, the condition $\nabla^{c}R\nabla_{c}R=0$ is equivalent to the standard definition of apparent horizon in terms of the expansion scalars of ingoing and outgoing null geodesic congruences [35, 28]. (if they exist) is $\nabla^{c}R\nabla_{c}R=0$. A single root corresponds to a black hole or white hole horizon, a double root to a wormhole [28]. This equation locates also event horizons, and gives
$g^{tt}\dot{R}^{2}+g^{rr}R^{\prime 2}=-\frac{\dot{R}^{2}}{T^{2}}+\frac{R^{\prime 2}}{a^{2}}=0$ (2.3)
where $\dot{R}\equiv\partial R/\partial t$, $R^{\prime}\equiv\partial R/\partial r$. Equation (2.3) is obtained also by imposing that a spherical surface of radius $R$ be null: $N_{a}\equiv\nabla_{a}R$ with $N^{a}N_{a}=0$. Indeed, it is well known that static apparent horizons coincide with null event horizons.
At a fixed instant of time $t=$ const., the event horizon is a 2-dimensional surface, a 2-sphere $R=$ const. spanned by the angular coordinates $\left(\vartheta,\varphi\right)$. The two-dimensional metric induced on the surface $R=$ const. is static but the time-dependent geometry just outside the horizon has $T=T(t,r)$, $a=a(t,r)$ and the areal radius $R(t,r)=a(t,r)r$ varies in time and cannot match $R=$ const. on the would-be horizon. Therefore, the junction conditions [36, 37] cannot be satisfied there: the discontinuity in the metric $g_{ab}$ induced on this 2-surface cannot be handled with distributional techniques and the jump in the metric $g_{ab}$, unlike a jump in its transversal derivative (interpreted as due to a matter layer on a null shell) has no physical interpretation [37].
III Radial geodesics and Ricci scalar at the horizon
Let us examine outgoing and ingoing radial null and timelike geodesics. We will show that they cannot be continued inside the would-be static black hole horizon, which is unphysical. Furthermore, the Ricci scalar of the geometry (2.1) diverges there. The difficulties arise from assuming an exactly static horizon in a time-dependent geometry, showing that this assumption is unphysical.
III.1 Radial null geodesics
Consider radial null geodesics parametrized by an affine parameter $\lambda$ and with 4-tangent $u^{a}=\left(u^{0},u^{1},0,0\right)$. The normalization $u_{c}u^{c}=0$ reads
$-T^{2}(u^{0})^{2}+a^{2}(u^{1})^{2}=0\,,$ (3.1)
yielding the coordinate velocity of light rays
$\frac{u^{1}}{u^{0}}=\frac{dr}{dt}=\pm\frac{T}{a}\,,$ (3.2)
where the upper sign refers to outgoing, and the lower one to ingoing, radial null geodesics. The equation of these affinely parameterized geodesics is
$\frac{du^{a}}{d\lambda}+\Gamma^{a}_{bc}u^{b}u^{c}=0\,.$ (3.3)
Using (3.2), the time component gives
$\frac{du^{0}}{d\lambda}+\left(\Gamma^{0}_{00}\pm\frac{2T}{a}\,\Gamma^{0}_{01}+\frac{T^{2}}{a^{2}}\,\Gamma^{0}_{11}\right)(u^{0})^{2}=0$ (3.4)
which, using the Christoffel symbols of the metric (2.1)
$\Gamma^{0}_{00}=\frac{\dot{T}}{T}\,,\quad\Gamma^{0}_{01}=\frac{T^{\prime}}{T}\,,\quad\Gamma^{0}_{11}=\frac{a\dot{a}}{T^{2}}\,,$ (3.5)
assumes the form
$\frac{d}{d\lambda}\left(\frac{1}{u^{0}}\right)=\frac{\dot{T}}{T}\pm\frac{2T^{\prime}}{a}+\frac{\dot{a}}{a}\,.$ (3.6)
Similarly, using the Christoffel symbols
$\Gamma^{1}_{00}=\frac{TT^{\prime}}{a^{2}}\,,\quad\Gamma^{1}_{01}=\frac{\dot{a}}{a}\,,\quad\Gamma^{1}_{11}=\frac{a^{\prime}}{a}\,,$ (3.7)
the radial component of the geodesic equation (3.3) gives
$\frac{d}{d\lambda}\left(\frac{1}{u^{1}}\right)=\frac{T^{\prime}}{T}\pm\frac{2\dot{a}}{T}+\frac{a^{\prime}}{a}\,.$ (3.8)
Close to the static event horizon, the spherical geometry must resemble Schwarzschild which, in isotropic coordinates, is well known to be approximated by the Rindler metric
$ds^{2}=ds^{2}_{\mathrm{Rindler}}+\,...=-\frac{x^{2}}{4}dt^{2}+4m^{2}\left(dx^{2}+d\Omega_{(2)}^{2}\right)+\,...\,,$ (3.9)
where $m$ is the black hole mass and $x$, given by $r=\frac{m}{2}\left(1+x\right)$ (with $0<x\ll 1$), measures the radial distance to the black hole horizon [38]. The horizon geometry is exactly static, hence $\dot{T}=\dot{a}=0$ there and Eq. (3.6) approximates to
$\frac{d}{d\lambda}\left(\frac{1}{u^{0}}\right)\simeq\pm\frac{2T^{\prime}}{a}\simeq\pm\frac{2}{a_{\mathrm{H}}}$ (3.10)
since $T^{\prime}\simeq 1$, which yields $u^{0}\simeq\frac{C}{\lambda-\lambda_{0}}$, with $C$ and $\lambda_{0}$ constants chosen so that $u^{0}$ is positive to keep the 4-tangent future-oriented. Then, either $u^{0}\to 0$ as $\lambda\to+\infty$, i.e., “time stops” (if $\lambda>\lambda_{0}$ and $C>0$) or else the radial null geodesic stops at the finite value $\lambda_{0}$ of $\lambda$ (if $\lambda<\lambda_{0}$ and $C<0$), with $u^{0}$ still diverging. On the contrary, $u^{1}=\pm Tu^{0}/a\to 0$ and the coordinate velocity (3.2) tends to zero as $x\to 0^{+}$, i.e., the radial motion of light stops at the horizon (for ingoing radial null geodesics) or cannot begin there (for outgoing ones). The fact that ingoing radial photons cannot penetrate the static event horizon and fall into the black hole is unphysical. The study of radial timelike geodesics gives similar results.
III.2 Radial timelike geodesics
We now turn to timelike geodesics. The 4-velocity $u^{a}=dx^{a}/d\tau$ (where $\tau$ is the proper time) tangent to a radial timelike geodesic has components $\left(u^{0},u^{1},0,0\right)$ and is normalized to $u^{c}u_{c}=-1$, which gives
$T^{2}\left(u^{0}\right)^{2}=a^{2}\left(u^{1}\right)^{2}+1\,.$ (3.11)
Since $(u^{0})^{2}=\frac{1+(au^{1})^{2}}{T^{2}}\geq\frac{1}{T^{2}}\to+\infty$ as $x\to 0$, $u^{0}$ diverges as least as fast as $T^{-1}=2/x$ at the static horizon. The time and radial components of the geodesic equation are (using $\dot{a},T\to 0$)
$\frac{du^{0}}{d\tau}\pm\frac{2u^{0}}{a}\,\sqrt{T^{2}(u^{0})^{2}-1}\,\frac{T^{\prime}}{T}=0\,,$ (3.12)
$\frac{du^{1}}{d\tau}\pm\frac{1+a^{2}(u^{1})^{2}}{a^{2}}\,\frac{T^{\prime}}{T}=\frac{a^{\prime}}{a^{3}}\,,$ (3.13)
where, in (3.13), we used the normalization (3.11) $u^{0}=\sqrt{1+\left(au^{1}\right)^{2}}/T$. Near the horizon, Eq. (3.13) becomes
$\frac{du^{1}/d\tau}{(u^{1})^{2}+1/a^{2}}=-\frac{1}{T}\,,$ (3.14)
which can be integrated at the horizon where $T$ is static, obtaining
$u^{1}\simeq\frac{1}{a}\tan\left(\frac{\tau_{0}-\tau}{aT}\right)\,,$ (3.15)
where $\tau_{0}$ is an integration constant. As $x\to 0$, $u^{1}$ is only defined for a finite range of the proper time $\tau$. Radially infalling objects stop at the horizon and cannot enter it.
III.3 Ricci scalar
Geodesic incompleteness is the signature of spacetime singularities in the Hawking-Penrose singularity theorems [39, 40], signaling the breakdown of spacetime at the static horizon. The other common definition of spacetime singularity, the divergence of curvature scalars, confirms this conclusion.
The Ricci scalar of the static geometry (2.1) (not to be confused with the areal radius $R$) is
${\cal R}$ $=$ $-\frac{4a^{\prime\prime}}{a^{3}}+\frac{6\ddot{a}}{aT^{2}}-\frac{2a^{\prime}T^{\prime}}{a^{3}T}-\frac{6\dot{a}\dot{T}}{aT^{3}}+\frac{6\dot{a}^{2}}{a^{2}T^{2}}-\frac{8a^{\prime}}{a^{3}r}$ (3.16)
$-\frac{4T^{\prime}}{a^{2}rT}-\frac{2T^{\prime\prime}}{a^{2}T}+\frac{2a^{\prime 2}}{a^{4}}\,.$
Because the horizon is exactly static, it must be $\dot{a}=0$ and $\dot{T}=0$ in the limit $x\to 0$, where (3.16) reduces to
${\cal R}\simeq-\frac{4a^{\prime\prime}}{a^{3}}-\frac{2a^{\prime}T^{\prime}}{a^{3}T}-\frac{8a^{\prime}}{a^{3}r}-\frac{4T^{\prime}}{a^{2}rT}-\frac{2T^{\prime\prime}}{a^{2}T}+\frac{2a^{\prime 2}}{a^{4}}\,.$ (3.17)
The first and the last terms in the right-hand side of Eq. (3.17) are finite and can be neglected in comparison with the terms diverging as $1/T$; using $T\sim x/2$, $T^{\prime}\simeq 1$, $T^{\prime\prime}\simeq 0$ and denoting with $r_{\mathrm{H}}$ and $a_{\mathrm{H}}$ the finite values of $r$ and $a$ on the horizon, one obtains
${\cal R}\simeq-\frac{2}{a_{\mathrm{H}}^{2}}\left[\left(\frac{a_{\mathrm{H}}^{\prime}}{a_{\mathrm{H}}}+\frac{2}{r_{\mathrm{H}}}\right)\frac{T^{\prime}}{T}+\frac{T^{\prime\prime}}{a_{\mathrm{H}}T}\right]\simeq\frac{1}{x}\to\infty$ (3.18)
as $x\to 0^{+}$. Similar considerations hold for the Kretschmann and other curvature scalars.
IV Test Klein-Gordon field
Matter is also subject to divergences at the would-be static horizon. Let us consider a free massless test scalar field $\phi$ satisfying the Klein-Gordon equation
$\Box\phi=\frac{1}{\sqrt{-g}}\,\partial_{a}\left(\sqrt{-g}\,g^{ab}\partial_{b}\phi\right)=0\,,$ (4.1)
where $g=\mbox{Det}(g_{ab})$. We are interested in the near-horizon propagation of the $\phi$-waves and, for simplicity, we restrict to $s$-modes with $\phi=\phi(t,r)$. In the geometry (2.1), Eq. (4.1) reads
$\ddot{\phi}+\left(\frac{3\dot{a}}{a}+\frac{\dot{T}}{T}\right)\dot{\phi}-\frac{T^{2}}{a^{2}}\,\phi^{\prime\prime}-\frac{T}{a^{2}}\left(T^{\prime}+\frac{Ta^{\prime}}{a}+\frac{2T}{r}\right)\phi^{\prime}=0\,.$ (4.2)
Since the would-be horizon is exactly static, $\dot{a}=0$, $\dot{T}=0$ and $T\to 0$, at the horizon the $s$-waves satisfy $\ddot{\phi}\simeq 0$ and $\phi(t)=\phi_{0}+\phi_{1}t$ (where $\phi_{0,1}$ are constants), diverging as $t\to+\infty$. Their energy-momentum tensor $T_{ab}=\nabla_{a}\phi\nabla_{b}\phi-\frac{1}{2}\,g_{ab}\nabla^{c}\phi\nabla_{c}\phi$ has trace
${T^{a}}_{a}=\left(\frac{\dot{\phi}}{T}\right)^{2}\to+\infty\quad\quad\mbox{as }\;x\to 0^{+}\,.$ (4.3)
The free scalar field is equivalent to an effective perfect fluid with stiff equation of state, effective energy density $\rho$, effective pressure $P$, and trace $T=-\rho+3P$, which diverges at the static horizon. Since $\rho=T_{00}=\dot{\phi}^{2}/2$ is finite, the effective pressure diverges at the horizon.
V Hawking temperature
Assume now that the time-dependent geometry is asymptotically FLRW, describing an exactly static black hole horizon embedded in a FLRW universe. The Hawking temperature at the horizon is $T_{\mathrm{H}}=1/(8\pi m)$, where $m$ is the black hole mass, taken as the Misner-Sharp-Hernandez mass defined in spherical symmetry (which reduces to the Schwarzschild mass) and must be constant for the black hole horizon to remain static. The black hole must be in equilibrium with its Hawking radiation, but the temperature of a photon gas in local thermal equilibrium when decoupled from other cosmic fluids scales as $T\sim 1/a$ to maintain the blackbody distribution. To wit, a physical wavelength scales as $\lambda_{\mathrm{phys}}=\lambda_{\mathrm{c}}a$, where $\lambda_{\mathrm{c}}$ is the comoving wavelength, the frequency $\nu=c/\lambda_{\mathrm{phys}}\sim 1/a$ and the Planck distribution of the spectral energy density
$u(\nu,T)=\frac{8\pi h\nu^{3}}{c^{3}}\,\frac{1}{\mbox{e}^{\frac{h\nu}{K_{B}T}}-1}$ (5.1)
(where $h$ and $K_{B}$ are the Planck and the Boltzmann constants, respectively) is only preserved if $T\sim 1/a$.
Furthermore, the analysis of exact solutions of the Einstein equations representing black holes embedded in FLRW universes, which are usually described by dynamical apparent (but not event) horizons, yields black hole temperatures scaling as $T\sim\frac{1}{8\pi ma}$. This is the case for cosmological black hole solutions of the Einstein equations obtained by conformally transforming Schwarzschild [41], in particular for the Sultana-Dyer black hole scrutinized in greater detail [42]. At our would-be static horizon it must be simultaneously $T=1/(8\pi m)=$ const. and $T\sim 1/a(t)$, which is impossible.
VI Conclusions
An exactly static black hole event horizon embedded in the time-dependent geometry (2.1) is, in reality, a naked null spacetime singularity located at a finite areal radius.
All radial null and timelike geodesics are future-inextendible at this static horizon. [^4] The Hawking-Penrose theorems identify a spacetime singularity where at least one geodesic is future-inextendible [39, 40]. Geodesic incompleteness defines spacetime singularities in the Hawking-Penrose singularity theorems [39, 40]. The Ricci scalar also diverges at the would-be horizon. A radially propagating test scalar field has divergent effective pressure there. Finally, the Hawking temperature of a strictly static black hole horizon would have to be constant, while the expansion of the universe in which it is embedded dictates that it must scale like the inverse scale factor, which are incompatible requirements.
The analysis in [34] assumed the Einstein equations and a perfect fluid with linear barotropic equation of state $P=\mbox{const.}\times\rho$ as a source. Here, we significantly generalized that result and we did not use any of these assumptions. Furthermore, we did not assume that the geometry is asymptotically FLRW (except in Sec. V) but only that it is time-dependent outside the static event horizon, therefore our result is quite general, including FLRW as a special case.
Forcing an exactly static black hole event horizon onto a dynamic “background” (e.g., a FLRW universe) turns this horizon into a naked null singularity. Preventing it from expanding (which can only be done with strong gravity objects since weak-field objects will comove) tears spacetime. The problems described and the unavoidability of the naked null singularity arise from the assumption that the black hole horizon is exactly static and disappear if this assumption is dropped, thus time-dependent event horizons are possible. Finding their dependence on time or on the scale factor of the host FLRW universe presumably requires the complete specification of the geometry and committment to specific metrics, and will be investigated in the near future.
The prospects of embedding an exactly stationary black hole event horizon in a FLRW (or any time-dependent) “background” are even more dire than for a spherical static horizon. In fact, a FLRW universe is not stationary and matching a horizon with axisymmetry (or any symmetry) with another metric that does not share this symmetry is doomed to failure. This fact is well known when trying to generalize the Swiss-cheese model (a Schwarzschild region matched to spherically symmetric FLRW on a 2-sphere) to axisymmetry. Matching a Kerr region to FLRW fails because the interior and exterior regions do not share axisymmetry [43, 44, 45]. This consideration, which will be discussed further in the future, extends the physical relevance of the result presented here.
Finally, we have not determined the precise way in which black hole event horizons embedded in given FLRW universes evolve with time. This law would be most interesting for astrophysics to decide whether existing puzzles can be solved by cosmological coupling, and will be addressed in the future.
Acknowledgements.
We are grateful to Kevin Croker and Duncan Farrah for discussions. V.F. is supported by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234).